# Dated Data: Tracing Knowledge Cutoffs in Large Language Models

### v1

@@ 4 Results
In this section we discuss the results of probing language models to determine their effective cutoffs and compare the determined dates with the ground truth distributions from their training sets.
Figure 3: Relative perplexities of models per month of the NewsSpan (NYT) dataset (§3.1). Relative perplexities are shown as the absolute perplexities do not matter for determining resource-level effective knowledge cutoffs. We find that our approach identifies the effective knowledge cutoff as the stated knowledge cutoff for NYT, as models have increased perplexity when their CommnonCrawl data dumps end in early 2020.
@@ 4.2 WikiSpan
As discussed in Section 3.3, there are three major categories of datasets that models are trained on: C4, the Pile, and Falcon RefinedWeb (RW); here we show results for each category. Again note that results show relative perplexities as absolute perplexities between different models are not relevant to our goal of determining knowledge-cutoffs. For each category, we also overlay the computed distribution of similar ground truth documents in light grey (when available).
Figure 4: Relative perplexities of models per month using the WikiSpan3.1) dataset (we use relative as exact perplexities are not needed for determining effective cutoffs). Upper plot shows Pile derived models, middle shows FalconRW derived models, while lower shows C4 derived models. The light grey bars indicate the distribution of Wikipedia-alike documents, matched to their closest version, as calculated in Section 3.5. Note that these datasets are only a subset of the training set for some models. In some cases the knowledge cutoff aligns with the model’s effective cutoff (e.g. the Pile) while for more recent models they are aligned much earlier (e.g. RedPajamas to 2019, even though it has an explicit 2023 Wikipedia dump).
@@
The three models GPT-Neo, GPT-J and Pythia are exclusively trained on the Pile. We show the results of the perplexity measurements in Fig. 4 (upper), where we see a noticeable drop in perplexity around March 2020. This month corresponds exactly to the date of the Wikipedia included in the Pile, indicating the effective cutoff for Wikipedia of the Pile-based models aligns with their reported Wikipedia cutoff. Appropriately, we also note that the distribution of ground truth versions is highest at that month, corresponding to the 3x up-sampled Wikipedia dump in the Pile.
@@
FalconRW is exclusively trained on RW while Falcon incorporates curated corpora from the Pile. Note that both models were trained on a subset of the RW dataset, the exact subset which is not publicly described. We see in Fig. 4 (middle) that the perplexity curves have a low perplexity from late 2019 to early 2021 with a minimum around January 2020. These results may seem surprising as FalconRW has not seen Wikipedia in training – however, as the overlayed ground truth and Section 5.1 shows, it does contain Wikipedia from other top-level domains.
@@
Fig. 4 (lower) shows results for C4-derived models: RedPajamas, OLMo, and LLaMa. While each model uses C4 during pre-training, it only comprises a small portion of their respective training data. The more salient similarity is that each of the models consists of many independent CommonCrawl dumps, and their differences in effective cutoff dates of the three models can be explained by the CommonCrawl dumps included in their training data. LLaMA incorporates 5 dumps from 2017 to 2020, and thus is the lowest during that time range. In contrast, RedPajamas incorporates 5 dumps from 2019 to 2023, and thus sees a minimum perplexity at a later date. OLMo uses all 20 dumps from 2019 to 2023, and thus sees a minimum perplexity at the latest date. Nonetheless, we still see the effect of C4 on these models, as the respective minima are biased towards the C4 CommonCrawl dump date of April 2019. See Section 5.2 for an explicit breakdown of the entire training data of RedPajamas.
4.3 Impact of Scale
We also consider the effect of model size on our methods. We evaluate the perplexity of WikiSpan under a suite of Pythia (460M, 1B, 2.8B, 6.9B, 12B) and LLaMA (7B, 13B, 65B) models. Fig. 5 shows that while the smaller models have a more varying perplexity curve, they are still minimized at the expected date. This makes intuitive sense as the models are all trained on the same data.
@@ 5 Why are models not aligned to their cutoff date?
In this section we describe why a model’s effective cutoff and reported cutoff can differ. This misalignment is due to two main factors: (1) deduplication pipelines that ignore semantically equivalent but lexically near duplicates and (2) the older time alignment of CommonCrawl dumps.
@@ 5.1 Complications in Deduplication Pipelines
It is common practice in LLM training pipelines to deduplicate data. When a dataset undergoes fuzzy or exact deduplication, one might expect that different versions of Wikipedia pages (near duplicates) and copies of Wikipedia pages (exact duplicates), respectively, are removed from the dataset. However, we empirically find that there exist a large number of near and exact duplicates in training datasets, and provide examples for each.
@@
FalconRW removed all documents that had the top-level domain of Wikipedia, so they could use FalconRW in conjunction with a curated version of Wikipedia in the future (as they did in the unreleased main Falcon dataset). They assumed this would deduplicate the data, however, we find that due to versioning and Wikipedia mirrors, there are still near duplicate Wikipedia documents, as shown in Table 2, where the texts differ only by the reference numbers.
Pile
Similarly, the Pile also contains accidentally duplicated Wikipedia documents. However, due to the explicit 3x Wikipedia up-sampling, these accidental duplicates are relatively minor in comparison.
@@
C4 was created from one CommonCrawl dump and “discarded all but one of any three-sentence span occurring more than once." However, we show an example in Table 3 of a pair of documents which contain the same three-sentence span. Beyond this, we can also see that the documents are semantically equivalent, and differ only by whitespace characters.
@@
We find that the RedPajamas CommonCrawl dump, which underwent paragraph-level deduplication, contains exact duplicate documents. We show an example in Table 4.
@@ Discussion
We confirm this hypothesis by comparing Pythia vs. Pythia-deduplicated. Fig. 6 shows that the deduplicated Pythia, which removes the purposefully up-sampled Wikipedia documents, no longer has the sharp minimum of standard Pythia and instead has a earlier effective cutoff (due to the older CC documents). Thus, we see that the accident duplicates and the lack of purposeful duplicates (of versions corresponding to the desired effective knowledge cutoff) creates this misalignment.
@@
Figure 7: Distribution of Wikipedia versions over the entire RedPajamas training set. Each color represents a different supcorpora of the training set (green for CommonCrawl, orange for Wikipedia, and blue for C4). The black line represents the perplexity curve of RedPajamas 7B over WikiSpan
@@ 5.2 Misalignment of CommonCrawl Dump Dates
All our evaluated models are trained on some portion of CommonCrawl data, with recent models using large proportions of it. Our ground truth results in Fig. 4 (especially C4-derived models) confirms previous work from Dodge et al. (2021) that suggest that a non-trivial amount of data inside of a CommonCrawl dump is actually old data. In the context of Wikipedia, this means that a CommonCrawl dump in 2023 often contains many versions of documents dating back to 2016. Thus, even though models may include CommonCrawl dumps from a range of dates, the aggregated data will be extremely biased towards earlier dates.
To illustrate this concretely for a “newer" style model composed of large amounts of CommonCrawl, we collect the ground truth from all the resources containing Wikipedia in RedPajamas (CommonCrawl, C4, and explicit Wikipedia) and overlay its perplexity curve in Fig. 7.777Note that although we performed this analysis by binning documents by their most similar version, one can compute an n-gram anaylsis with similar results (Appendix C).
@@ 5.3 Summary
Thus, we see that there exists two reasons that contribute to the temporal misalignment of language model’s reported cutoff vs. the effective cutoff: (1) failures of current deduplication pipelines to control for semantic duplicates and (2) the usage and reliance on newer CommonCrawl dumps to provided update information when in actuality there is a significant amount of older data.
@@ 6 Conclusion
It is now common practice for Large Language Models to provide a “knowledge-cutoff" which intends to communicate to users the date at which LLMs no longer have up to date information. However, this simple metric oversimplifies LLM training in a detrimental manner to usability; it leaves unanswered the questions of “is this knowledge cutoff specific for all resources in the model", “how many copies of my resource are in the model" or “which versions of my corpus are included?" We propose a method to automatically determine the effective cutoff date of LLMs for a given resource and show that although sometimes it does align with the reported cutoff, in many cases it does not. To determine why they fail to align, we analyze the training data of open-data LLMs to discover that there are large quantities of near-duplicates in LLM training data (such as differing only in the citation numbers included in the text) despite efforts from LLM creators to deduplicate. Further, most LLMs rely on CommonCrawl dumps for data, despite the fact that a non-trivial amount of CommonCrawl data is much older than the reported dump date. We hope this analysis will provide insight for users of LLMs who need resource-specific knowledge cutoffs and for LLM-creators who seek to align their LLMs to a given date.
@@ Appendix A Probing Pseudocode
We denote the pre-training dataset $\mathcal{D}$. Let $*M*$ denote the month corresponding to the reported Wikipedia dump in $\mathcal{D}$. Let $D_{T,M}$ represent the version of topic $T$ at month $M$. edit refers to the Levenshtein distance between two strings.
@@ Appendix B Deduplication Complications
"Adam Richard Sandler (born September 9, 1966) is an American actor, comedian, screenwriter, film producer, and musician. After becoming a Saturday Night Live cast member, he went on to star in many Hollywood feature films that have grossed over \$2 billion at the box office combined.Sandler’s well-known roles include Billy Madison (1995), Happy Gilmore (1996), The Waterboy (1998), The Wedding Singer (1998), Big Daddy (1999), Mr. Deeds (2002), 50 First Dates (2004), The Longest Yard (2005), Click (2006), Grown Ups (2010), Just Go with It (2011), Grown Ups 2 (2013), Blended (2014), and Murder Mystery (2019). He also voices Dracula in the Hotel Transylvania franchise (2012–present). Some of his films, such as the widely panned Jack and Jill, have been heavily criticized, culminating in a shared second place in the number of Raspberry Awards (3) and Raspberry Award nominations (11), in both cases second only to Sylvester Stallone. Sandler ventured into dramatic territory with his roles in Punch-Drunk Love (2002), Spanglish (2004), Reign Over Me (2007), Funny People (2009), The Meyerowitz Stories (New and Selected) (2017), and Uncut Gems (2019), all of which earned him critical praise."
@@ Appendix C N-gram Analysis instead of Exact Match
How is perplexity affected when seeing two lexically similar texts? One hypothesis is that perplexity is most affected by exact and near-duplicates. Another hypothesis is that the actual text in a document is factor affecting perplexity. To illustrate the difference between these hypotheses, a document that contains shuffled sentences from a version of a Wikipedia article (shuffling order of sentences) would not affect perplexity in the former case, but would in the latter case. The formeh hypothesis is the basis behind our algorithm for attributing matched documents to their closest versions. We test the latter hypothesis by proposing another way to attribute credit, directly counting the intersection of $n$-grams in matched documents with precomputed sets of $n$-grams sourced from WikiSpan. We discount by the number of times an ngram appears across all months (similar to an inverse document frequency) in order to count ngrams that are distinct to a specific Wikipedia version. Our algorithm and results are described in Section C.1

### v2

@@ 5 Results
In this section we discuss the results of probing language models to determine their effective cutoffs and compare the dates with the ground truth distributions from their training sets.
@@ 5.1 NewsSpan
Figure 3: Relative perplexities of models per month using the NewsSpan3.1) dataset (we use relative as exact perplexities are not needed for determining effective cutoffs). We find that our approach identifies the effective cutoffs as the stated knowledge cutoff for NYT, as models have increased perplexity when their CommnonCrawl data dumps end in 2020.
@@ 5.2 WikiSpan
As discussed in Section 4.1, there are three major categories of datasets that models are trained on: C4, the Pile, and Falcon RefinedWeb (RW) and we note that the datasets are only a subset of the training set for some models. Again only uncomparable relative perplexities are shown as absolute perplexities between different models are not relevant to our goal of determining knowledge-cutoffs. For each category, we also overlay the computed distribution of similar ground truth documents in light grey (when available) to show the correlation between the ground truth results and effective cutoffs.
@@
The three models GPT-Neo, GPT-J and Pythia are exclusively trained on the Pile. We show the results of the perplexity measurements in Fig. 4 (upper), where we see a noticeable drop in perplexity around March 2020. This month corresponds exactly to the date of the Wikipedia included in the Pile, indicating the effective cutoff for Wikipedia of the models aligns with the reported Wikipedia cutoff. Appropriately, we also note that the distribution of ground truth versions is highest at that month, corresponding to the 3x up-sampled Wikipedia dump in the Pile.
@@
FalconRW is exclusively trained on RW while Falcon incorporates curated corpora from the Pile. Note that both models were trained on a subset of the RW dataset, the exact subset which is not publicly described. We see in Fig. 4 (middle) that the perplexity curves have a low perplexity from late 2019 to early 2021 with a minimum around January 2020. These results may seem surprising as FalconRW has not seen Wikipedia in training – however, as the overlayed ground truth and Section 6.1 shows, it does contain Wikipedia content found on other top-level domains.
@@
Fig. 4 (lower) shows results for C4-derived models: RedPajamas, OLMo, and LLaMa. While each model uses C4 during pre-training, it only comprises a small portion of their respective training data. The more salient similarity is that each of the models consists of many independent CommonCrawl dumps, and the differences in effective cutoff dates of the three models can be explained by the CommonCrawl dumps included in their training data. LLaMA incorporates 5 dumps from 2017 to 2020, and its cutoff date is thus in that range. In contrast, RedPajamas incorporates 5 dumps from 2019 to 2023, and its effective cutoff is a few months later. OLMo uses all 20 dumps from 2019 to 2023, and thus sees the latest effective cutoff. Nonetheless, the effect of C4 on these models is evidenced by the effective cutoffs being biased towards the C4 CommonCrawl dump date (April 2019). See Section 6.2 for a breakdown of the entire training data of RedPajamas.
Figure 4: Relative perplexities of models per month using the WikiSpan3.1) dataset. Upper plot shows Pile derived models, middle shows FalconRW derived models, while lower shows C4 derived models. The light grey bars indicate the distribution of Wikipedia-alike documents, matched to their closest version, as calculated in Section 3.3. In some cases the knowledge cutoff aligns with the model’s effective cutoff (e.g. the Pile) while more recent models are aligned much earlier (e.g. RedPajamas to 2019, even though it has an explicit 2023 Wikipedia dump).
Impact of Scale
We also consider the effect of model size on our methods. We evaluate the perplexity of WikiSpan under a suite of Pythia (460M, 1B, 2.8B, 6.9B, 12B) and LLaMA (7B, 13B, 65B) models. Fig. 5 shows that while the smaller models have a more varied perplexity curve, they are still minimized at the expected date. This makes intuitive sense as the models are all trained on the same data.
@@
Number of Documents
We lastly consider the effect of the number of documents measured. In some domains, document collection may be difficult; as such, we evaluate our method by varying $x$, the number of documents considered (2500, 1000, 500, 250, 100, 50) for the three C4 derived models. Fig. 6 shows for $x>50$, the effective cutoffs of the three models are consistent with the full results. $x=50$ appears to be the threshold where the trend are less consistent, likely due to the increased variance in the data. This shows that our method is robust even when many versions of documents cannot be collected.
Figure 6: Relative perplexities of RedPajamas (top), LLaMA (middle), and OLMo (bottom) when varying $x$, the number of documents in each bucket. Darker colors indicate more documents, and the black lines corresponding to $x=5000$ are the results shown in Fig. 4. For small $x$, the perplexity curves are more variable due to the smaller sample size, but for $x>50$, the ablated results are consistent with the full results.
@@ 6 Why are models not aligned to their cutoff date?
In this section we describe why a model’s effective cutoff and reported cutoff can differ. This mismatch is due to two main factors: (1) deduplication pipelines that ignore semantically equivalent but lexically near duplicates and (2) temporal biases of CommonCrawl dumps.
@@ 6.1 Complications in Deduplication Pipelines
It is common practice in LLM training pipelines to deduplicate data. In the context of Wikipedia, when a dataset undergoes fuzzy or exact deduplication, one might expect that different versions of Wikipedia pages (near duplicates) and copies of Wikipedia pages (exact duplicates), respectively, are removed from the dataset. However, we empirically find that there exist a large number of near and exact duplicates in training datasets, and provide examples for each. This shows that deduplication pipelines are unable to detect the extra copies and versions of Wikipedia documents present in CommonCrawl dumps.
@@
FalconRW removed all documents that had Wikipedia as a top level domain so they could use FalconRW in conjunction with curated versions in the future (as they did in for the Falcon dataset). They assumed this would deduplicate the data, however, we find that there are still near duplicate Wikipedia documents, as shown in Table 2.
@@
C4 was created from one CommonCrawl dump and “discarded all but one of any three-sentence span occurring more than once.” However, we show an example in Table 3 of a pair of documents which contain the same three-sentence span. We also observe that the documents are semantically equivalent, and differ only by whitespace characters.
@@
We find that the RedPajamas CommonCrawl dump that was paragraph-level deduplicated contains exact duplicate documents. We show an example in Table 4.
@@ Discussion
We confirm this hypothesis by comparing Pythia vs. Pythia-deduplicated. Fig. 7 shows that the deduplicated Pythia, which removes the purposefully up-sampled Wikipedia documents, no longer has the sharp minimum of standard Pythia and instead has an earlier effective cutoff (due to the older CC documents). Thus, we see that the accidental duplicates and the lack of purposeful duplicates (of versions corresponding to the desired effective knowledge cutoff) creates this misalignment in the deduped models.
@@
Figure 8: Distribution of Wikipedia versions over the entire RedPajamas training set. Each color represents a different supcorpora of the training set. The black line represents the relative perplexity curve of RedPajamas 7B over WikiSpan.
@@ 6.2 Misalignment of CommonCrawl Dump Dates
All our evaluated models were trained on some portion of CommonCrawl data, with recent models using larger proportions of it. Our ground truth results in Fig. 4 (especially C4-derived models) confirm previous work from Dodge et al. (2021) that suggests a non-trivial amount of data inside of a CommonCrawl dump is actually old data. In the context of Wikipedia, this means that a CommonCrawl dump in 2023 contains many versions of documents dating back to 2016. While models may include a range of CommonCrawl dumps, the aggregated data will thus be extremely biased towards earlier dates. To illustrate this concretely for a “newer” style model composed of large amounts of CommonCrawl, we collect the ground truth from all the resources containing Wikipedia in RedPajamas (CommonCrawl, C4, and explicit Wikipedia) and overlay its perplexity curve in Fig. 8.777Note that although we performed this analysis by binning documents by their most similar version, one can compute an n-gram anaylsis with similar results (Appendix C).
@@ 6.3 Summary
In our analysis of available pretraining datasets, we find that CommonCrawl dumps often include multiple copies of different versions of documents (e.g. Wikipedia). These extra copies and versions are frequently undetected by deduplication pipelines and moreover can consist of outdated information, biasing the effective cutoffs of language models. Thus, we see that there exists two reasons that contribute to the temporal mismatch of a language model’s reported and effective cutoff: (1) failures of deduplication pipelines to control for semantic duplicates and (2) the use of newer CommonCrawl dumps to provide updated information when there is a significant amount of older data in the dumps.
@@ 7 Conclusion
It is now common practice for Large Language Models to provide a “knowledge-cutoff” which intends to communicate to users the date at which LLMs no longer have up to date information. However, this simple metric oversimplifies LLM training in a detrimental manner to usability; it leaves unanswered the questions of “is this knowledge cutoff specific for all resources in the model”, “how many copies of my resource are in the model” or “which versions of my corpus are included?” We propose a method to automatically determine the effective cutoff date of LLMs for a given resource and show that although sometimes it does align with the reported cutoff, in many cases it does not. To determine why they fail to align, we analyze the training data of open-data LLMs to discover that there are large quantities of near-duplicates in LLM training data (such as differing only in the citation numbers included in the text) despite efforts from LLM creators to deduplicate. Further, most LLMs rely on CommonCrawl dumps for data, despite the fact that a non-trivial amount of CommonCrawl data is much older than the reported dump date. We hope this analysis will provide insight for users of LLMs who need resource-specific knowledge cutoffs and for LLM-creators who seek to align their LLMs to a given date.
8 Acknowledgements
We thank JHU HLTCOE for the support and resources used for this project. We additionally thank the students in JHU CLSP for the valuable feedback, particularly Nathaniel Weir, Kate Sanders, Zhengping Jiang, Jingyu (Jack) Zhang and Aleem Khan. OW is supported by the National Science Foundation Graduate Research Fellowship Program. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Probing Pseudocode
We denote the pre-training dataset $\mathcal{D}$. Let $M*$ denote the month corresponding to the reported Wikipedia dump in $\mathcal{D}$. Let $D_{T,M}$ represent the version of topic $T$ at month $M$. edit refers to the Levenshtein distance between two strings.
@@ B.1 Falcon
B.3 RedPajamas
”Adam Richard Sandler (born September 9, 1966) is an American actor, comedian, screenwriter, film producer, and musician. After becoming a Saturday Night Live cast member, he went on to star in many Hollywood feature films that have grossed over \$2 billion at the box office combined.Sandler’s well-known roles include Billy Madison (1995), Happy Gilmore (1996), The Waterboy (1998), The Wedding Singer (1998), Big Daddy (1999), Mr. Deeds (2002), 50 First Dates (2004), The Longest Yard (2005), Click (2006), Grown Ups (2010), Just Go with It (2011), Grown Ups 2 (2013), Blended (2014), and Murder Mystery (2019). He also voices Dracula in the Hotel Transylvania franchise (2012–present). Some of his films, such as the widely panned Jack and Jill, have been heavily criticized, culminating in a shared second place in the number of Raspberry Awards (3) and Raspberry Award nominations (11), in both cases second only to Sylvester Stallone. Sandler ventured into dramatic territory with his roles in Punch-Drunk Love (2002), Spanglish (2004), Reign Over Me (2007), Funny People (2009), The Meyerowitz Stories (New and Selected) (2017), and Uncut Gems (2019), all of which earned him critical praise.”
@@ Appendix C N-gram Analysis instead of Exact Match
How is perplexity affected when seeing two lexically similar texts? One hypothesis is that perplexity is most affected by exact and near-duplicates. Another hypothesis is that the actual text in a document is factor affecting perplexity. To illustrate the difference between these hypotheses, a document that contains shuffled sentences from a version of a Wikipedia article (shuffling order of sentences) would not affect perplexity in the former case, but would in the latter case. The former hypothesis is the basis behind our algorithm for attributing matched documents to their closest versions. We test the latter hypothesis by proposing another way to attribute credit, directly counting the intersection of $n$-grams in matched documents with precomputed sets of $n$-grams sourced from WikiSpan. We discount by the number of times an ngram appears across all months (similar to an inverse document frequency) in order to count ngrams that are distinct to a specific Wikipedia version. Our algorithm and results are described in Section C.1